# When Worse is Better: Navigating the compression-generation tradeoff in visual tokenization

### v1

@@ 3.4 Sequence Length and Compute Scaling
In this section, we investigate how scaling the number of tokens per image affects the rFID and gFID.
Setup. To vary the number of tokens per image, we follow [43] to re-use our tokenizers trained at $256\times 256$ resolution and evaluate them at higher resolutions. Our base tokenizers trained use a downsample factor of 16: if we process an $R\times R$ image, the number of tokens produced is $(R/16)^{2}=R^{2}/256$. To measure rFID and PSNR, we then take the $R\times R$ output and downsize to $256\times 256$ with bicubic interpolation. In Figure 5 we observe the expected result that increasing sequence length improves reconstruction performance. We opt for this method as opposed to changing the number of downsample layers in the stage 1 tokenizer (e.g. [9, 61, 27]) to avoid introducing a confounding variables related to architectural changes.
Model Capacity and Sequence Length. In Figure 4, we study how performance scaling changes as a result of both sequence length and stage 2 model size. We vary the training iterations at three model scales,  [^1] M, 340M, and 550M parameters and study how the number of tokens per image affects generation performance. We see that for smaller models, training with a lower number of tokens per image is more compute efficient, whereas larger models can take advantage of the greater representation capacity of longer sequence lengths. The exact training parameters and full pareto curves for these experiments can be found in the appendix. Notably in those results, we find that increasing sequence lengths makes the compute optimal Pareto curve worse above the 256 token/image rFID lower bound.
Figure 6: Codebook reconstruction performance scaling. We observe that reconstruction performance reliably improves as we increase codebook size across all metrics. We compare vector quantization (VQ, orange) with finite scalar quantization (FSQ, blue), a popular recent “lookup-free” method. In contrast to prior work, we find that with a good codebook recipe, VQ uniformly outperforms FSQ and scales to $>$100k codes.
Figure 7: Pareto-optimal image-generation compute curves for different codebook sizes: We plot the pareto-optimal compute curves for stage 2 performance for different codebook sizes (1k, 16k, and 131k). In the low-compute regime ($10^{15}$ FLOPs), we observe that the 1k codebook outperforms the 131k by 5-10 gFID; in the high compute regime ($10^{18}$ FLOPs), we observe the reverse effect. Notably, the 16k codebook consistently matches or outperforms 1k and 131k at every compute scale, indicating a sweet spot for codebook size. Full scaling plots are in the appendix.
Figure 8: Stage 2 compute scaling laws with CRT tokens (ImageNet 256x256). We demonstrate the superior scaling of CRT tokens across four orders of magnitude with respect to validation cross entropy (left) and FID (center). The trade-off between reconstruction performance and causal dependence means that CRT has worse reconstruction (2.36 rFID vs 2.21 for the baseline) but yields better stage 2 performance across many model scales. The result is a tokenizer that has better absolute generation performance for per parameter scale, and gets the same performance with 1.5×1.5$\times$1.5 × to 2.5×2.5$\times$2.5 × fewer training FLOPs (when comparing pareto frontiers).
@@ 3.5 Codebook Size and Compute Scaling
Codebook size is the second mechanism for changing the rate of the image tokenizer. This differs from changing sequence length because it has a minor effect on inference cost, both for the stage 1 and stage 2 models. Changing sequence length increases inference cost quadratically in stage 2 and linearly in stage 1, so the effect is more subtle.
Stage 1 scaling by codebook size. In Figure 6, we find consistent improvements in stage 1 rFID, PSNR, and MS-SSIM without collapse in performance, in contrast to prior work [9, 32, 55]. We credit this scaling improvement to modern VQ recipes [53, 43] in conjunction with a long learning rate warm-up period (training recipe details can be found in the appendix). Notably, the gap for rFID between the extremes of our exploration (1k and 131k) is 1.3 rFID.
Comparison with finite-scalar quantization [32]. In Figure 6, compare VQ to FSQ, which is an increasingly popular “lookup-free” quantization scheme across the recommended settings. We find that with a good VQ training recipe, FSQ is outperformed at all codebook scales in both stage 1 and stage 2, in contrast to the results demonstrated in [25, 32]. As a result, we primarily use VQ in this work.
Stage 2 model scale and codebook size. In Figure 7, we plot the Pareto frontiers for the extremes of the codebook sizes we study (1k and 131k), along with our baseline codebook size of 16k. We only show the Pareto frontiers for clarity but plot the full model scaling curves in the appendix. When we use less training compute, we see that the 1k codebook outperforms 131k codebook by 5-10 gFID, even though the rFID is significantly lower. At the other extreme, 131k outperforms 1k, as the stage 2 model nears the gFID saturation point and the worse rFID becomes a bottleneck. This crossover happens around $10^{17}$ train FLOPs. It’s worth noting that the 16k codebook size outperforms both at almost every compute scale, implying that there is an optimal point for the rFID/gFID trade-off.
@@ 3.6 Causally Regularized Tokenization
We showed that stage 2 model capacity is important when designing a stage 1 tokenizer. Going past capacity, since we know our stage 2 model is an auto-regressive transformer, can we imbue this inductive bias into the tokenizer?
Setup. Figure 8 demonstrates our key intervention. We apply an $\ell_{2}$ loss parameterized by a causal transformer to the tokens before quantization, and backpropagate this loss to the encoder of the our stage 1 auto-encoder. We train with this model-based loss ($\mathcal{L}_{CRT}$) in conjunction standard VQGAN losses, using the same hyperparameters as our baseline recipe. Applying this loss with cross-entropy after quantization is intuitive, however we found this unstable since VQ tokens change drastically during training. We use a 2-layer causal transformer with the architecture of our stage 2 model. This introduces a 5% cost in training FLOPs, which we control for by training 5% less (2 fewer epochs). For both our baseline and CRT models we use a codebook of size 16k and 256 tokens per image.
Effects on stage 1 reconstruction. After applying this loss, our tokenizer is worse in terms of reconstruction against the compute normalized baseline (Figure 5). This is expected behavior. Intuitively, our loss attempts to make token $i$ as predictable as possible given tokens 0 through $i-1$. The most predictable set of tokens would be if the encoder outputs a sequence of entirely dependent tokens (e.g., the constant token). However, this reduces the total information that can be conveyed through the tokens. Our CRT loss (a proxy for generation) conflicts with reconstruction, implying a trade-off. With this inductive bias, stage 2 scaling is greatly improved, especially at small to medium scales.
Improvements in stage 2 performance. In Figure 8, we show.the relative scaling properties of stage 2 models using our CRT tokenizer compared to the VQ-GAN baseline (in Figure 3). For clarity, we also include Pareto frontier scaling trends from Figure 3. Figure 8 (left) Shows that stage 2 models trained with our tokenizer achieve lower validation losses. Figure 8 (center) shows that this translates to improvements generation performance on the training compute Pareto front. Finally, Figure 8 (right) shows the actual magnitude of improvement, which the log-log-linear plots obfuscate: CRT achieves gFID scores with 1.53×1.5-3$\times$1.5 - 3 × compute efficiency and improve over the baseline, even with worse rFID. We demonstrate non-cherrypicked qualitative outputs from these models in the appendix. We show comparisons at individual model scales in Table 2, and detailed comparisons across training durations in the appendix.
LSUN [52]. We demonstrate our method’s generality by evaluating several LSUN categories. We do not reproduce the full scaling study due to computational cost, but in Table 3 we demonstrate improvements in gFID and CLIP [36] gFID across three model scales using CRT. Note that we use CLIP gFID because these datasets are out of distribution for the Inception network, making FID a more noisy metric. We see a consistent improvement across all model scales.
Table 1: System-level comparison between methods on ImageNet 256x256: We compile results from previous work that report stage 2 autoregressive image generation scores. Our tokenizer method (CRT) generally outperforms other methods when matching number of parameters of the image generation model. CRTopt refers to our hyperparameter-tuned CRT tokenizer.
@@
Diff. ADM [6] 554M 10.94 101.0 0.69 0.63
Diff. LDM-4-G [40] 400M 3.60 247.7
Diff. DiT-L/2 [33] 458M 5.02 167.2 0.75 0.57
Diff. DiT-XL/2 [33] 675M 2.27 278.2 0.83 0.57
Diff. MAR-H [30] 943M 1.55 303.7 0.81 0.62
Mask. MaskGIT [3] 227M 256 6.18 182.1 0.80 0.51
Mask. RCG (cond.) [29] 502M 256 3.49 215.5
VAR VAR-$d$16 [45] 310M 680 3.30 274.4 0.84 0.51
VAR VAR-$d$20 [45] 600M 680 2.57 302.6 0.83 0.56
VAR VAR-$d$24 [45] 1.0B 680 2.09 312.9 0.82 0.59
VAR VAR-$d$30 [45] 2.0B 680 1.92 323.1 0.82 0.59
AR VQVAE-2 [37] 13.5B 5120 31.11 $\sim$45 0.36 0.57
AR VQGAN [9] 227M 256 18.65 80.4 0.78 0.26
AR VQGAN [9] 1.4B 256 15.78 74.3
AR VQGAN-re [9] 1.4B 256 5.20 280.3
AR ViTVQ [53] 1.7B 1024 4.17 175.1
AR RQTran. [27] 3.8B 256 7.55 134.0
AR LlamaGen-B [43]  [^1] M 256 6.10 182.54 0.85 0.42
AR LlamaGen-L [43] 343M 256 3.80 248.28 0.83 0.51
AR LlamaGen-XL [43] 775M 256 3.39 227.08 0.81 0.54
AR LlamaGen-XXL [43] 1.4B 256 3.09 253.61 0.83 0.53
AR LlamaGen-B [43]  [^1] M 576 5.46 193.61 0.83 0.45
AR LlamaGen-L [43] 343M 576 3.07 256.06 0.83 0.52
AR LlamaGen-XL [43] 775M 576 2.62 244.08 0.80 0.57
AR LlamaGen-XXL [43] 1.4B 576 2.34 253.90 0.80 0.59
AR LlamaGen-3B [43] 3.1B 576 2.18 263.33 0.81 0.58
@@
Table 2: Compute-controlled comparison against tuned VQ-GAN baseline on ImageNet: We show image generation performance when we control all compute factors (train and test FLOPs for both stage 1 and stage 2 models, with stage 2 trained to saturation, 16k codebook size). Across stage 2 model sizes we observe that CRT is generally superior, indicating that tokens with causal-regularization are more amenable to stage 2 auto-regressive training. See Figure 8 for a comprehensive sweep.
@@
CRT (ours) (2.36 rFID) 740M 256 2.35 259.12 0.81 0.59
Table 3: Compute-controlled comparison against tuned VQ-GAN across different datasets: We provide compute-controlled comparison against our tuned VQGAN baseline (16k codebook size) for different training sets. Similar to Table 2, we observe that CRT generally outperforms a tuned VQ-GAN baseline.
@@
LSUN Cats Baseline (1.35 rFID)  [^1] M 256 5.77 8.14 0.58 0.56
340M 256 5.21 7.37 0.58 0.58
740M 256 5.31 6.72 0.59 0.59
 [^1] M 256 5.32 7.22 0.60 0.54
340M 256 4.98 6.77 0.61 0.56
CRT (ours) (1.48 rFID) 740M 256 5.32 6.74 0.60 0.57
LSUN Horses  [^1] M 256 3.19 5.56 0.62 0.62
340M 256 3.00 5.09 0.62 0.62
Baseline (1.65 rFID) 740M 256 2.56 4.92 0.62 0.63
 [^1] M 256 2.65 5.20 0.61 0.61
340M 256 2.62 4.63 0.61 0.63
CRT (ours) (1.83 rFID) 740M 256 2.76 4.05 0.61 0.63
LSUN Bedrooms Baseline (0.99 rFID)  [^1] M 256 2.57 9.81 0.59 0.53
340M 256 2.51 9.61 0.58 0.55
740M 256 2.53 9.50 0.57 0.57
 [^1] M 256 2.37 9.06 0.60 0.53
340M 256 2.08 8.56 0.59 0.56
CRT (ours) (1.39 rFID) 740M 256 2.22 8.50 0.58 0.57
@@
Optimizing for absolute performance. In the previous sections, we perform compute-controlled comparisons of CRT and baseline. In this section, we apply simple optimizations from literature to our tokenizer to push the limits of our method. These include: (1) increasing training duration from 38 epochs to 80 epochs, (2) doubling the size of the decoder, (3) increasing the codebook size from 16k to 131k (which improves performance for CRT but not our baseline). With these changes, our tokenizer achieves 1.21 rFID. Our stage 2 training procedures remain constant. In Table 1, we show a system-level comparison of our method (CRTopt-AR) to other existing methods in the field. The closest comparison to our method is LlamaGen, which uses the same architectures and inference method. Notably, we exceed the performance of the LlamaGen 3.1B parameter model with our 775M parameter model, while using half the number of tokens per image. This is an 8x reduction in inference compute. We show non-cherry-picked qualitative examples from select object categories in Figure 2.
@@
Ablations. There are two key choices involved with our causal loss: (1) the size of the causal model used during stage 1 training and (2) the weight $\lambda$ we assign to the $\ell_{2}$ next-token prediction loss ($\lambda\mathcal{L}_{CRT}$). We include an ablation of these parameters in Figure 9. We plot the relationship of rFID and gFID with respect to these parameters, computing gFID over an average of 4 stage 2 model scales (80M,  [^1] M, 211M, 340M) trained for 250k iterations each. In Figure 9 (top), we increase the number of layers in our causal regularizer. For every 2-layers we add, we reduce the overall training time by 5% to compensate. We see that rFID uniformly increases with more layers, as expected. After 4 layers, this worsening rFID hurts average gFID. Thus, we chose to use two layers in all of our experiments, which has the added advantage of being very computationally cheap. In Figure 9 (bottom), we see a similar trend. Increasing $\lambda$ biases the encoder to focus on making representations easy for next-token prediction rather than reconstruction, making overall rFID worse. However, there’s a sweet spot around $\lambda=4$ where this worsened reconstruction is within acceptable range and this also improves overall generation performance.

### v2

@@ 3.4 Sequence Length and Compute Scaling
In this section, we investigate how scaling the tokens per image affects rFID and gFID.
Setup. To vary the number of tokens per image, we follow [sun2024autoregressive] to re-use our tokenizers trained at $256\times 256$ resolution and evaluate them at higher resolutions. Our base tokenizers are trained with a downsample factor of 16: if we process an $R\times R$ image, the number of tokens produced is $(R/16)^{2}=R^{2}/256$. To measure rFID and PSNR, we take the $R\times R$ output and downsize to $256\times 256$ with bicubic interpolation. In Figure 3 we observe the expected result that increasing sequence length improves reconstruction performance. It also outperforms increasing codebook size, however at the expense of a linear increase in inference cost. We use this method as opposed to changing the number of encoder downsample layers (e.g. [esser2020taming, zheng2022movq, lee2022autoregressive]) to avoid introducing confounding variables related to architecture and to provide more granular control over token counts.
Sequence Length and Scaling Laws. In Figure 5, we see that for smaller models, training with a lower number of tokens per image is more compute efficient, whereas larger models take advantage of the greater representation capacity of longer sequence lengths. Notably in those results, we find that increasing sequence lengths makes the compute optimal Pareto curve better only close to the 256 token/image rFID lower bound. Our main conclusion is that more compressed sequences are generally more compute optimal before performance saturation. As gFID approaches the rFID barrier, using more tokens per image becomes compute optimal.
@@ 3.5 Codebook Size and Compute Scaling
Codebook size is the second main mechanism for changing the rate of the image tokenizer. This differs from changing sequence length because it has a minor effect on inference cost, both for the stage 1 and stage 2 models, so the effect on scaling is more subtle. This is in contrast to scaling sequence length, which increases inference cost linearly on both stage 2 and stage 1.
Stage 1 scaling by codebook size. In Figure 3, we show consistent improvements from 1k to 131k codebook size in stage 1 rFID, PSNR, and MS-SSIM without collapse in performance or codebook utilization, in contrast to prior work [esser2020taming, mentzer2023finite, yu2023language]. We credit this scaling improvement to modern VQ recipes [yu2021vector, sun2024autoregressive] in conjunction with a long learning rate warm-up period (training recipe details can be found in the appendix). We also show that our VQ recipe outperforms finite-scalar quantization (FSQ) [mentzer2023finite], a modern popular non-VQ quantization scheme.
Stage 2 model scale and codebook size. In Figure 3, we study scaling laws for the extremes of the codebook sizes we study (1k and 131k), along with our baseline codebook size of 16k. When we use less training compute, we see that the 1k codebook outperforms 131k codebook by 5-10 gFID, even though the rFID is significantly lower. At the other extreme, 131k outperforms 1k, as the stage 2 model nears the gFID saturation point and the worse rFID becomes a bottleneck. This crossover happens around $10^{17}$ train FLOPs. Important to note is that the 16k codebook size outperforms both at almost every compute scale, implying that there is an optimal point for the rFID/gFID trade-off. Interestingly, the change in scaling is not nearly as extreme as that exhibited by the tokens per image ablation in Section 3.4, despite the massive difference in rFID between 1k, 16k and 131k (1.7vs 2.21 vs 3.0 rFID). We suggest that this is because changing codebook size leaves training and inference FLOPs largely unaffected, while compute scales linearly with increased token counts. We discuss this further in Section 4. This also notably differs with prior works [chen2112simple, yu2023language], which show major differences between stage 2 performance when using VQ with large codebook sizes.
@@ 3.6 Causally Regularized Tokenization
We have shown how Pareto optimal scaling laws change based on bits per pixel (bpp) used for tokenization. We have also shown that increasing bits per pixel above 16k codebook size and 256 tokens per image generally makes scaling worse. Further, while reducing to 1k codebook size improves the scaling law marginally, explicitly reducing the reconstruction capacity to such a degree greatly harms generation performance close to saturation. Going past bpp as a measure of rate, we study an alternate rate-distortion trade-off based on optimizing stage 1 latents for stage 2 performance. Our guiding question is: since our stage 2 model is an auto-regressive transformer, can we imbue this inductive bias into the tokenizer? We show we can take our optimal baseline tokenizer and apply a simple causal regularization that improves scaling without changing bpp used for compression.
Figure 7: Stage 2 compute scaling laws with CRT tokens (ImageNet 256x256). We demonstrate the superior scaling of CRT tokens across four orders of magnitude with respect to validation cross entropy (left) and FID (center). The trade-off between reconstruction performance and causal dependence means that CRT yields better stage 2 performance across many model scales with worse reconstruction (2.36 vs 2.21 rFID). The result is improved absolute generation performance and $1.5\times$ to $2.5\times$ better scaling by training FLOPs.
Table 1: System-level and compute-controlled comparisons of autoregressive image-generation models. (a) System-level comparison on ImageNet 256×256. (b) Compute-controlled comparison against a tuned VQ-GAN baseline on ImageNet. (c) Compute-controlled comparison against a tuned VQ-GAN baseline across LSUN Cats, Horses, and Bedrooms. Across all settings our tokenizer (CRT) and its reconstruction performance optimized variant, CRTopt, outperform baselines.
@@
[…]
Diff. ADM [dhariwal2021diffusion] 554M 10.94 101.0 0.69 0.63
Diff. LDM-4-G [rombach2021high] 400M 3.60 247.7
Diff. DiT-L/2 [peebles2022scalable] 458M 5.02 167.2 0.75 0.57
Diff. DiT-XL/2 [peebles2022scalable] 675M 2.27 278.2 0.83 0.57
Diff. MAR-H [li2024autoregressive] 943M 1.55 303.7 0.81 0.62
Mask. MaskGIT [chang2022maskgit] 227M 256 6.18 182.1 0.80 0.51
Mask. RCG (cond.) [li2024return] 502M 256 3.49 215.5
VAR VAR-dd16 [tian2024visual] 310M 680 3.30 274.4 0.84 0.51
VAR VAR-dd20 [tian2024visual] 600M 680 2.57 302.6 0.83 0.56
VAR VAR-dd24 [tian2024visual] 1.0B 680 2.09 312.9 0.82 0.59
VAR VAR-dd30 [tian2024visual] 2.0B 680 1.92 323.1 0.82 0.59
AR VQVAE-2 [razavi2019generating] 13.5B 5120 31.11 $\sim$45 0.36 0.57
AR VQGAN [esser2020taming] 227M 256 18.65 80.4 0.78 0.26
AR VQGAN [esser2020taming] 1.4B 256 15.78 74.3
AR ViTVQ [yu2021vector] 1.7B 1024 4.17 175.1
AR LlamaGen-B [sun2024autoregressive]  [^1] M 256 6.10 182.54 0.85 0.42
AR LlamaGen-L [sun2024autoregressive] 343M 256 3.80 248.28 0.83 0.51
AR LlamaGen-XL [sun2024autoregressive] 775M 256 3.39 227.08 0.81 0.54
AR LlamaGen-XXL [sun2024autoregressive] 1.4B 256 3.09 253.61 0.83 0.53
AR LlamaGen-B [sun2024autoregressive]  [^1] M 576 5.46 193.61 0.83 0.45
AR LlamaGen-L [sun2024autoregressive] 343M 576 3.07 256.06 0.83 0.52
AR LlamaGen-XL [sun2024autoregressive] 775M 576 2.62 244.08 0.80 0.57
AR LlamaGen-XXL [sun2024autoregressive] 1.4B 576 2.34 253.90 0.80 0.59
AR LlamaGen-3B [sun2024autoregressive] 3.1B 576 2.18 263.33 0.81 0.58
@@
(b) Comparison against tuned VQGAN baseline (ImageNet) at saturation. CRT uniformly out-performs the baseline
@@
CRT (ours) (2.36 rFID) 775M 256 2.35 259.12 0.81 0.59
(c) Comparison against tuned VQ-GAN across LSUN datasets: Under the same compute constraints as (b), CRT uniformly surpasses the tuned baseline on LSUN Cats, Horses, and Bedrooms. gFIDCLIP is the primary metric because Inception features are out-of-distribution for LSUN.
@@
LSUN Cats Baseline (1.35 rFID)  [^1] M 256 8.14 5.77 0.58 0.56
340M 256 7.37 5.21 0.58 0.58
775M 256 6.72 5.31 0.59 0.59
 [^1] M 256 7.22 5.32 0.60 0.54
340M 256 6.77 4.98 0.61 0.56
CRT (ours) (1.48 rFID) 775M 256 6.74 5.32 0.60 0.57
LSUN Horses  [^1] M 256 5.56 3.19 0.62 0.62
340M 256 5.09 3.00 0.62 0.62
Baseline (1.65 rFID) 775M 256 4.92 2.56 0.62 0.63
 [^1] M 256 5.20 2.65 0.61 0.61
340M 256 4.63 2.62 0.61 0.63
CRT (ours) (1.83 rFID) 775M 256 4.05 2.76 0.61 0.63
LSUN Bedrooms Baseline (0.99 rFID)  [^1] M 256 9.81 2.57 0.59 0.53
340M 256 9.61 2.51 0.58 0.55
775M 256 9.50 2.53 0.57 0.57
 [^1] M 256 9.06 2.37 0.60 0.53
340M 256 8.56 2.08 0.59 0.56
CRT (ours) (1.39 rFID) 775M 256 8.50 2.22 0.58 0.57
Setup. Figure 1 demonstrates our key intervention. We apply an $\ell_{2}$ loss parameterized by a causal transformer to the tokens before quantization, and backpropagate this loss to the encoder of the our stage 1 auto-encoder. We train with this model-based loss ($\mathcal{L}_{CRT}$) in conjunction standard VQGAN losses, using the same hyperparameters as our baseline recipe. Applying this loss with cross-entropy after quantization is intuitive; however, we found that this damages reconstruction fidelity too much (see Section 4), since VQ tokens change discretely during training. If token $i$ changes to token $j$, the cross-entropy loss treats it the same regardless of how similar the tokens $i$ and $j$ are. Meanwhile, $\ell_{2}$ on pre-quantized tokens are naturally similarity aware. We use a 2-layer causal transformer with the architecture of our stage 2 model. This introduces a 5% cost in training FLOPs, which we control for by training 5% less (2 fewer epochs). For both our baseline and CRT models we use a codebook of size 16k and 256 tokens per image.
Effects on stage 1 reconstruction. After applying this loss, our tokenizer is marginally worse in terms of reconstruction against the compute normalized baseline (2.36 rFID vs 2.21 rFID). Intuitively, our loss attempts to make token $i$ as predictable as possible given tokens 0 through i1i-1. The most predictable set of tokens would be if the encoder outputs a sequence of entirely dependent tokens (e.g., the constant token). However, this reduces the total information that can be conveyed through the tokens. Our CRT loss (a proxy for generation) conflicts with reconstruction, implying a trade-off. With this inductive bias, stage 2 scaling is improved, which we discuss next.
Improvements in stage 2 performance. In Figure 7, we show the relative scaling properties of stage 2 models using our CRT tokenizer compared to the VQ-GAN baseline (in Figure 4). For clarity, we also include Pareto frontier scaling trends from Figure 4. Figure 7 (left) Shows that stage 2 models trained with our tokenizer achieve systematically lower validation losses. Figure 7 (center) shows that this translates to improvements generation performance on the training compute Pareto front. Namely, our compute scaling law is improved from $\alpha=-0.65$ to $\alpha=-0.73$. Finally, Figure 7 (right) shows the actual magnitude of improvement, which the log-log-linear plots obfuscate: CRT achieves gFID scores with $1.5-3\times$ compute efficiency and improve over the baseline, even with worse rFID. We show comparisons at individual model scales in Table 1b, and detailed comparisons across training durations in Appendix F. Our model systematically outperforms the baseline in all compute and model size regimes.
Results on LSUN [yu2015lsun]. In Table 1c, demonstrate our method’s generality by evaluating several LSUN categories. We demonstrate improvements in gFID and gFIDCLIP across three model scales using CRT across all model scales. Note that we use CLIP gFID because these datasets are out of distribution for the Inception network, making FID a more noisy metric.
@@
Optimizing the CRT tokenizer for absolute performance. In the previous sections, we perform compute-controlled comparisons of CRT and baseline. CRT outperforms the baseline despite the trade-off in reconstruction performance. However, we showed previously that generation performance close to the rFID barrier is affected by reconstruction quality. In this section, we apply simple optimizations to boost reconstruction performance, which do not modify our fundamental setup, to push the limits of our method. These include: (1) increasing training duration from 38 epochs to 80 epochs, (2) doubling the size of the decoder, (3) increasing the codebook size from 16k to 131k, which interestingly improves performance for CRT but not our baseline. With these changes, our tokenizer achieves 1.26 rFID. Our stage 2 training procedures remain constant.
In Table 1a, we show a system-level comparison of our method (CRTopt-AR) to other existing methods in the field. The closest comparison to our method is LlamaGen, which uses the same architectures and inference method. Notably, we exceed the performance of the LlamaGen 3.1B parameter model with our 775M parameter model, while using half the number of tokens per image. This is an 8x reduction in inference compute. We show uncurated qualitative examples from select object categories in Figure 14 and further examples in Appendix G.
@@
Hyperparameter Ablations. There are two key choices involved with our causal loss: (1) the size of the causal model used during stage 1 training and (2) the weight $\lambda$ we assign to the $\ell_{2}$ next-token prediction loss ($\lambda\mathcal{L}_{CRT}$). We include an ablation of these parameters in Figure 8. We plot the relationship of rFID and gFID with respect to these parameters, computing gFID over an average of 4 stage 2 model scales (52M,  [^1] M, 211M, 340M) trained for 250k iterations each. In Figure 8 (left), we increase the number of layers in our causal regularizer. For every 2-layers we add, we reduce the overall training time by 5% to compensate. We see that rFID uniformly increases with more layers, as expected. After 4 layers, this worsening rFID hurts average gFID. Thus, we chose to use two layers in all of our experiments, which has the added advantage of being very computationally cheap. In Figure 8 (right), we see a similar trend. Increasing $\lambda$ biases the encoder to focus on making representations easy for next-token prediction rather than reconstruction, making overall rFID worse. However, there’s a sweet spot around $\lambda=4$ where this worsened reconstruction is within acceptable range and this also improves overall generation performance.
Loss Function Ablation ($\ell_{2}$ vs cross-entropy). In Appendix E, we show comparisons between $\ell_{2}$ loss and cross-entropy (CE) loss. We also compare to the loss suggested by [wang2024larptokenizingvideoslearned], who propose a similar auto-regressive inductive bias in the video domain (based on CE not $\ell_{2}$). In general, CE losses harm reconstruction performance in our setting but fail to provide substantial generation performance improvement to be worth the trade-off. This deviation from findings in [wang2024larptokenizingvideoslearned] is likely due to two reasons: (1) They use an architecture that is substantially different from ours, focusing on global versus local tokens (we use the VQGAN architecture for compatibility with existing SOTA literature) (2) The video generation setting they studied is much further from performance saturation, meaning changes to tokenizer reconstruction are less likely to harm generation performance.